# arXiv:2603.00918v1[cs.CV] 01 Mar 2026

### v1

@@ 10 Training Collapse Analysis
Mitigations used in ARC. We restrict training to the latter $60\%$ of steps ($\rho=0.6$), keep CFG on during rollouts (but off when scoring self-confidence), and retain clipping, per-timestep weighting, and antithetic probes. These choices suppress overconfidence spikes and stabilize learning.
Figure 7: Visualization of training collapse in ARC. Self-confidence (y-axis) versus training iteration under different settings. Using $\rho>0.6$ or sampling rollouts without CFG drives a steep, short-horizon increase in self-confidence, followed by degenerate outputs—evidence of reward hacking. ARC’s default settings ($\rho{=}0.6$ and CFG for rollouts) avoid this behavior while preserving steady improvements.
@@ 11 Additional Ablation Studies
We conduct additional ablation studies and comparative experiments to validate the design choices of ARC. The results are summarized in Tab. 5.
11.1 Caption datasets for ARC
ARC relies on intrinsic self-confidence and thus requires only prompts (not external reward models). We compare three prompt sources: (i) text-rendering (OCR) prompts from Flow-GRPO [37] (our default), (ii) PickScore [30] prompts, and (iii) GenEval [21] prompts. As shown in Tab. 3, denser, more prescriptive prompts (OCR) yield the strongest gains; empirically, self-confidence is most reliable when the text condition is explicit and descriptive. We provide the descriptions and examples for each prompt dataset in Tab. 4.
@@ 11.1 Caption datasets for ARC
Table 4: Prompt sources compared for ARC. Denser, text-focused prompts (OCR) provide stronger supervision signals for intrinsic self-confidence, leading to larger gains than more open-ended (PickScore) or simple compositional (GenEval) prompts.
@@ 11.2 Effect of group size
Although ARC’s self-confidence can be computed per step, we find that using the aggregated reward, i.e., averaging weighted per-step scores over the probed timesteps, consistently performs better than optimizing stepwise advantages. Stepwise improvements at individual timesteps need not translate to a better final sample and tend to increase variance and solver sensitivity; aggregation provides a more stable, outcome-aligned signal for post-training.
@@
Caption dataset used for ARC
@@ 12 Additional Implementation Details
In this section we summarize the main implementation choices used in our ARC training pipeline. We acknowledge and correct a typographical error in the main paper: although we stated that the group size was $G=24$, all experiments were in fact conducted with $G=16$. The summary of hyperparameterse and configurations is illustrated in Tab. 6.
Category Hyperparameter Value (ARC, SD3.5-M)
@@
Self-confidence (ARC) Probes per step $K$ 88 (antithetic pairing: $K/2$ noise, $K/2$ negated)
@@
Table 6: Hyperparameters and key implementation details for ARC training on SD3.5-M.
@@ 12.1 Base models and LoRA configuration
We build on the StableDiffusion3Pipeline from diffusers with the pretrained model SD3.5-M: stabilityai/stable-diffusion-3.5-medium. We freeze all components except the denoiser: the VAE and all text encoders are kept fixed and used only for inference. Only the main transformer (UNet-like denoiser) is updated during training, based on LoRA. We run the text encoders in mixed precision (fp16 in our main ARC runs) and keep the VAE in fp32 for stability.
@@ 14 Additional Qualitative Results
We provide side-by-side samples for (i) PickScore–post-trained (Flow-GRPO) SD3.5–M, (ii) FLUX.1–Dev, and (iii) SD3.5–L in Fig. 9. Across diverse prompts, ARC yields visibly sharper text rendering, more faithful object counts and relations, and fewer artifacts, echoing the quantitative gains in compositionality, text rendering, and text–image alignment, with no obvious regressions on non-target aspects.
Figure 9: Additional qualitative results of ARC. We present additional qualitative results of ARC when applied to (1) Flow-GRPO [37] post-trained SD3.5-M [16], (2) FLUX.1-Dev [3], and (3) SD3.5-L [16]. Best viewed on electronics.
Acknowledgement. This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale)

### v2

@@ 10 Training Collapse Analysis
Mitigations used in SOLACE. We restrict training to the latter $60\%$ of steps ($\rho=0.6$), keep CFG on during rollouts (but off when scoring self-confidence), and retain clipping, per-timestep weighting, and antithetic probes. These choices suppress overconfidence spikes and stabilize learning.
Figure 7: Visualization of training collapse in SOLACE. Self-confidence (y-axis) versus training iteration under different settings. Using $\rho>0.6$ or sampling rollouts without CFG drives a steep, short-horizon increase in self-confidence, followed by degenerate outputs—evidence of reward hacking. SOLACE’s default settings ($\rho{=}0.6$ and CFG for rollouts) avoid this behavior while preserving steady improvements.
@@ 11 Additional Ablation Studies
We conduct additional ablation studies and comparative experiments to validate the design choices of SOLACE. The results are summarized in Tab. 5.
11.1 Caption datasets for SOLACE
SOLACE relies on intrinsic self-confidence and thus requires only prompts (not external reward models). We compare three prompt sources: (i) text-rendering (OCR) prompts from Flow-GRPO [37] (our default), (ii) PickScore [30] prompts, and (iii) GenEval [21] prompts. As shown in Tab. 3, denser, more prescriptive prompts (OCR) yield the strongest gains; empirically, self-confidence is most reliable when the text condition is explicit and descriptive. We provide the descriptions and examples for each prompt dataset in Tab. 4.
@@ 11.1 Caption datasets for SOLACE
Table 4: Prompt sources compared for SOLACE. Denser, text-focused prompts (OCR) provide stronger supervision signals for intrinsic self-confidence, leading to larger gains than more open-ended (PickScore) or simple compositional (GenEval) prompts.
@@ 11.2 Effect of group size
Although SOLACE’s self-confidence can be computed per step, we find that using the aggregated reward, i.e., averaging weighted per-step scores over the probed timesteps, consistently performs better than optimizing stepwise advantages. Stepwise improvements at individual timesteps need not translate to a better final sample and tend to increase variance and solver sensitivity; aggregation provides a more stable, outcome-aligned signal for post-training.
@@
Caption dataset used for SOLACE
@@ 12 Additional Implementation Details
In this section we summarize the main implementation choices used in our SOLACE training pipeline. We acknowledge and correct a typographical error in the main paper: although we stated that the group size was $G=24$, all experiments were in fact conducted with $G=16$. The summary of hyperparameterse and configurations is illustrated in Tab. 6.
Category Hyperparameter Value (SOLACE, SD3.5-M)
@@
Self-confidence (SOLACE) Probes per step $K$ 88 (antithetic pairing: $K/2$ noise, $K/2$ negated)
@@
Table 6: Hyperparameters and key implementation details for SOLACE training on SD3.5-M.
@@ 12.1 Base models and LoRA configuration
We build on the StableDiffusion3Pipeline from diffusers with the pretrained model SD3.5-M: stabilityai/stable-diffusion-3.5-medium. We freeze all components except the denoiser: the VAE and all text encoders are kept fixed and used only for inference. Only the main transformer (UNet-like denoiser) is updated during training, based on LoRA. We run the text encoders in mixed precision (fp16 in our main SOLACE runs) and keep the VAE in fp32 for stability.
@@ 14 Additional Qualitative Results
We provide side-by-side samples for (i) PickScore–post-trained (Flow-GRPO) SD3.5–M, (ii) FLUX.1–Dev, and (iii) SD3.5–L in Fig. 9. Across diverse prompts, SOLACE yields visibly sharper text rendering, more faithful object counts and relations, and fewer artifacts, echoing the quantitative gains in compositionality, text rendering, and text–image alignment, with no obvious regressions on non-target aspects.
Figure 9: Additional qualitative results of SOLACE. We present additional qualitative results of SOLACE when applied to (1) Flow-GRPO [37] post-trained SD3.5-M [16], (2) FLUX.1-Dev [3], and (3) SD3.5-L [16]. Best viewed on electronics.